# LLM-Text Watermarking based on Lagrange Interpolation

### v1

@@ Abstract
The rapid advancement of LLMs (Large Language Models) has established them as a foundational technology for many AI- and ML-powered human–computer interactions. A critical challenge in this context is the attribution of LLM-generated text – either to the specific language model used or to the individual user who generated it. This is essential for combating misinformation, fake news, misinterpretation, and plagiarism. One of the key techniques for addressing this issue is watermarking.
This work presents a watermarking scheme for LLM-generated text based on Lagrange interpolation, which enables the recovery of a secret author identity even when the text has been heavily redacted by an adversary. The core idea is to embed a continuous sequence of points $(x,f(x))$ that lie on a single straight line. The $x$-coordinates are generated pseudorandomly using either an LFSR (when security is not a priority) or a cryptographically secure NFSR for high-security applications. The scheme’s efficiency and resilience to adversarial modifications are analysed. Experimental results show that the proposed method is highly effective, allowing the recovery of the author’s identity when as few as three points survive adversarial manipulation.
@@ 1 Introduction
The growing adoption of Large Language Models (LLMs) offers a wide array of practical benefits, including automated text correction, assistance with writing, and content refinement. However, alongside these advantages lies a more troubling aspect: the potential misuse of LLMs. One such concern arises when individuals present AI-generated content as their own, falsely claiming authorship. This challenge naturally raises an important question: how can we embed a watermark into LLM-generated text that allows a recipient to reliably trace its origin, whether to a specific user or to the LLM system itself?
The objective of such watermarking is to make it possible to attribute the text to its true source, thereby establishing authorship beyond reasonable doubt. This capability is especially critical in combating misinformation, propaganda, automated plagiarism, and the spread of fake news. In this paper, we take a step toward a simple yet effective watermarking method for LLM-generated text, aiming to create a foundation for trustworthy AI-generated communication.
@@ 2 Overview and Taxonomy of Watermarking in LLMs
– the output probability distribution (logits) is subtly perturbed before sampling tokens, embedding statistical patterns without altering the model weights. The work [12] splits the collection of tokens into green and red subsets. The green tokens encode watermarks while red ones mark fake ones. Wong et. al. [22] describes their logits-to-text watermarking, which jointly optimises encoder-decoder architecture for high-fidelity and high-detectability watermarks. The GumbelSoft watermarking described in [10] improves sampling diversity while maintaining robust detection by refining Gumbel-based logits control.
@@ 2.1 Contributions
    We design a watermarking scheme that embeds points from a straight line $f(x)$ into LLM-generated text. The $x$-coordinates are generated pseudorandomly, and the watermark consists of the sequence of points $(x,f(x))$. To extract the watermark, the verifier must solve the Maximum Collinear Points problem. We show that several algorithms exist for this task, with the most efficient ones having quadratic complexity, i.e., ${\cal O}(N^{2})$, where $N$ is the number of points.
  • $\bullet$
    We propose a variant of the watermarking scheme that is immune to algebraic attacks. This version employs cryptographically secure generation of $x$-coordinates and dynamic partitioning of the token set into ${\cal V}_{0}$ and ${\cal V}_{1}$.
@@
    The proposed scheme can be naturally extended to support multiple secrets. We present two options: (1) using multiple lines, each encoding a distinct secret, and (2) using higher-degree polynomials instead of lines. In the latter case, reconstructing the embedded polynomial corresponds to solving the Maximum Co-Polynomial Points problem. However, existing algorithms for this task have complexities that grow with the polynomial’s degree.
Table 2 compares some multi-bit watermarking schemes with the one presented in this work.
Table 1: Some LLM-generated text multi-bit watermarking schemes
Authors Tools User Identification Integrity Complexity / Efficiency
Qu et al [18] RS Codes, Hash RS decoding
Jiang et al [11] TTP, Hash Efficient – needs TTP
This work Lagrange Interpolation, Hash, NFSR Solving MCP
The rest of the paper is structured as follows. Section 3 focuses on the quality and security properties required from a watermarking scheme. Section 4 describes the tools and building blocks used in the design of our watermarking schemes. Section 5 introduces our basic watermarking scheme along with algorithms for recovering watermarking bits. Section 6 presents an algebraic analysis of the scheme and a variant that is immune to such attacks. It also includes a probabilistic analysis evaluating the likelihood that an adversary can implant a fake point that would be accepted as genuine. Section 7 explores techniques for fast watermark extraction and double-resolution watermarking. Section 8 details our experimental results. Section 9 demonstrates how our watermarking scheme can be extended to support multiple secrets. Finally, Section 10 concludes the paper.
@@ 3 Quality and Security Requirements for LLM-text Watermarking
In this work, we focus on a specific security goal: enabling the owner of an LLM-generated text to prove that the text contains a substantial portion of their original content, from which a secret identity $K$ can be reliably recovered. This identity serves as a cryptographic witness of authorship or ownership. We further show how to extend our scheme so it provides integrity verification. Our approach is grounded in the use of Lagrange interpolation over points in a two-dimensional plane. Each point encodes a fragment of the embedded identity. To ensure unpredictability and prevent pattern leakage, we employ a pseudorandom bit generator (PRBG) to determine the $x$-coordinates of the interpolation points.
@@ 4.1 Lagrange Interpolation over $GF(2^{n})$
4.2 Pseudorandom Bit Generation
We need to generate $x$-coordinates at random. To achieve this, we rely on generators that produce bit sequences exhibiting random-like behaviour, even though they are deterministic algorithms initialised by a seed. These generators generally fall into three main categories: linear feedback shift registers (LFSRs), nonlinear feedback shift registers (NFSRs), and cryptographically strong pseudorandom bit generators (CSPBGs).
A linear feedback shift register is a shift register with linear feedback, typically using XOR, that generates deterministic sequences from an initial state. With a well-chosen feedback polynomial, an $m$-bit LFSR can produce a maximum-length sequence (m-sequence) cycling through all $12^{m}-$ 1 nonzero states. LFSRs are valued for their simplicity, speed, and efficient hardware implementation, making them popular in cryptography, error detection (e.g., CRC), random number generation, and spread-spectrum communication. Despite their advantages, LFSRs are vulnerable to algebraic attacks – observing just $m$ output bits can reveal the full state, and even nonlinear filtering does not fully mitigate this risk (see Courtois and Meier [7]).
A nonlinear feedback shift register is a generalization of the LFSR, in which the feedback function is highly nonlinear. This added complexity increases resistance to cryptanalytic techniques such as linear and algebraic attacks. NFSRs are widely used in cryptographic applications due to their strong security properties and efficient implementation in both software and hardware. A notable selection of NFSR-based designs was proposed during the CAESAR competition [3].
A cryptographically strong pseudorandom bit generator is a deterministic algorithm that produces bit sequences that are computationally indistinguishable from true randomness. Unlike general-purpose pseudorandom generators, CSPBGs are specifically designed to resist cryptographic attacks, ensuring that an adversary cannot feasibly predict future output even after observing a polynomial number of bits. Their security is typically based on hard mathematical problems, such as the shortest vector problem in lattices. While highly secure, CSPBGs are often less efficient due to the computational complexity of the underlying cryptographic assumptions.
To maintain efficient implementation, LFSRs are preferred, particularly when the LLM-generated text is short, limiting the adversary’s ability to gather sufficient information to compromise the watermarking scheme, or when security is not the primary concern. In scenarios where stronger security is required, NFSRs serve as a robust alternative to LFSRs, offering enhanced resistance to cryptanalysis. CSPBGs, however, are generally not recommended for this context due to their inherent trade-off: achieving high security typically relies on solving computationally intractable problems, such as operations over high-dimensional lattices, which are resource-intensive and slow.
@@ 5 Watermark based on Lagrange Interpolation
In many cases, the owner of an LLM-generated text may need to prove authorship by demonstrating that the embedded watermark enables recovery of their secret identity $K$. In other words, the extraction algorithm should be capable of retrieving this identity even if the text has been modified by an adversary. To ensure verifiable authorship, any sufficiently long fragment of the watermarked text should contain enough information to reveal the owner’s identity, serving as undeniable proof. We propose Lagrange interpolation as a convenient tool for recovering a secret polynomial $f(x)$, whose coefficients encode the identity $K$, from points in the two-dimensional plane with randomly generated $x$-coordinates. This guarantees that any sufficiently large portion of the text yields enough points to reconstruct $K$. The approach is resilient to partial alterations, making it robust in adversarial scenarios where the text may be tampered with to hide authorship.
As previously mentioned, we use Lagrange interpolation for polynomials of degree $t=2$. In this case, we work with a straight line $f(x)=a_{0}+a_{1}x\in GF(2^{n})[x]$, where the secret identity is encoded as $K=(a_{0}\|a_{1})\in GF^{2}(2^{n})$. Suppose our watermark consists of $N$ points lying on $f(x)$, implying that the number of tokens in the LLM text is $T=2nN$. An LFSR with secret seed $s_{k}$ is used to generate the coordinates $x_{i}$. For each $x_{i}$, we compute the pair $(x_{i},f(x_{i}))$, resulting in the watermark sequence:
$(x_{1},f(x_{1})),\ldots,(x_{N},f(x_{N})),$
where $x_{i}=LFSR(t_{i},s_{k})$ and $t_{i}$ denotes the clock instance. Any two such points are sufficient to reconstruct the secret $K$. Now assume the original LLM text $\cal T$ has been tampered with by an adversary, resulting in a modified version $\widetilde{\cal T}$. A verification algorithm extracts the watermark from $\widetilde{\cal T}$, possibly containing errors or missing points. Crucially, all correctly recovered points will still lie on the original line $f(x)=a_{0}+a_{1}x$. The verifier can accept $K=(a_{0}\|a_{1})$ as a valid candidate if it corresponds to the line with the most points recovered from $\widetilde{\cal T}$. If a recovered point is $(\tilde{x}_{i},\widetilde{f(x_{i})})$ instead of the correct $(x_{i},f(x_{i}))$, the probability that it still lies on the correct line is $\nicefrac{{1}}{{2^{n}}}$. This holds because for any incorrect $\tilde{x}_{i}$, there exists exactly one value $f(\tilde{x}_{i})$ that would make the point valid.
Fortunately, our adversary is unlikely to remove, replace, or insert single tokens, as such changes are easily detectable as typos. In other words, a single change (that is not a typo) typically generates a sequence of changes enforced by the LLM statistics (or the adversary intention). For this reason, the adversary is more likely to operate on larger sequences of tokens (such as sentences or paragraphs). This also means that the adversary tampers with longer sequences of tokens. A typical English language sentence consists of 10 to 20 tokens, while a paragraph contains between 150 and 200 tokens. Moreover, there is a trend in the development of new LLMs that are capable of processing more and more tokens.
Algorithm 5 outlines the process of watermark embedding using Lagrange interpolation. We depart from the conventional approach where the lists ${\cal V}_{0},{\cal V}_{1}$ are selected anew for each token, as this introduces unnecessary computational overhead. Assume we wish to embed the bit $i\in\{0,1\}$. First, the LLM is queried to produce logits. We then adjust the probabilities of the top tokens (based on the logits) that belong to ${\cal V}_{i}$ by checking whether the corresponding parity condition is satisfied. The parity check may be defined either as the XOR of the token ID bits or, alternatively, as a parity function applied to the output bits of a hash function $H(s_{k},ID_{\mbox{\it token}})$ – see Section 4.3. In this approach, the algorithm uses static sets ${\cal V}_{0}$ and ${\cal V}_{1}$ for embedding watermark bits across the text. While it is straightforward to use dynamic lists to enhance security, doing so increases the computational cost of watermark extraction.
@@
Algorithm 1: Watermark Embedding Secret Identity $K$ 1:Input: Autoregressive language model LLM used to generate a text of length $T=2nN$, prompt sequence $PR$, token set ${\cal V}$, $n$-bit secret key $s_{k}$, LFSR with $n$-bit state and straight line $f(x)=a_{0}+a_{1}x$ with the secret identity $K=(a_{0}\|a_{1})$, where computations are done in $GF(2^{n})$. 2:Output: Public LLM text $S_{1,1},\dots,S_{N,2n}$.   3:Initialize $LFSR(t_{0},s_{k})$ 4:Using secret key $s_{k}$, create two lists ${\cal V}_{0}$, ${\cal V}_{1}$, where ${\cal V}_{i}$ are tokens carrying watermark $i\in\{0,1\}$ (${\cal V}_{0}\cup{\cal V}_{1}={\cal V}$) 5:for $i=1$ to $N$ do 6:     $x_{i}=(x_{i,1},x_{i,2},\ldots,x_{i,n})\leftarrow LFSR(t_{i},s_{k})$ 7:     $y_{i}=f(x_{i})=(y_{i,1},y_{i,2},\ldots,y_{i,n})$ 8:     for $j=1$ to $n$ do 9:         Compute logits: $v_{1}=\text{LLM}(PR,S_{1,1},\dots,S_{i,2(j-1)})$ 10:         Bias the logits: 11:         if $x_{i,j}=0$ then add bias to $\hat{v}_{1}[w]=v_{1}+\delta$ for all $w\in{\cal V}_{0}$ 12:         else add bias to $\hat{v}_{1}[w]=v_{1}+\delta$ for all $w\in{\cal V}_{1}$ 13:         end if 14:         Append $S_{i,2j-1}$ to $S$ by sampling from biased logits $\hat{v}_{1}$ 15:         Compute logits: $v_{2}=\text{LLM}(PR,S_{1,1},\dots,S_{i,2j-1})$ 16:         Bias the logits: 17:         if $y_{i,j}=0$ then add bias to $\hat{v}_{2}[w]=v_{2}+\delta$ for all $w\in{\cal V}_{0}$ 18:         else add bias to $\hat{v}_{2}[w]=v_{2}+\delta$ for all $w\in{\cal V}_{1}$ 19:         end if 20:         Append $S_{i,2j}$ to $S$ by sampling from biased logits $\hat{v}_{2}$ 21:     end for 22:end for 23:return $S$
@@
The algorithm processes text block by block. For each block (lines 6,7), two watermark sequences are prepared. The first is a block generated by LFSR, which consists of an $x$-coordinate. The second block $y_{i}=f(x_{i})$. The loop from line 8 to 21 injects bits of pairs $(x_{i},f(x_{i}))$ into the LLM text.
@@
Algorithm 2: Watermark Extraction 1:Input: LLM text $\widetilde{\cal T}=(\tilde{S}_{1,1},\dots,\tilde{S}_{N,2n})$, token set ${\cal V}$, $n$-bit secret key $s_{k}$, $n$-bit state LFSR, which generates $x$-coordinates of unknown $f(x)$, where points $(x_{i},f(x_{i}))$ are embedded into the text according to Algorithm 5; $i=1,\ldots,N$, computations are done in $GF(2^{n})$. 2:Output: The secret user identity $K=(a_{0}\|a_{1})$   3:Create an empty list of points ${\mathbb{P}}=\varnothing$ of $(x,f(x))$ 4:Initialize $LFSR(t_{0},s_{k})$ 5:Create the sequence ${\mathbb{L}}=(x_{1},\ldots,x_{N})$, where $x_{i}=LFSR(t_{i},s_{k})$ 6:Using secret key $s_{k}$, create two lists ${\cal V}_{0}$, ${\cal V}_{1}$, where ${\cal V}_{i}$ are tokens carrying watermark $i\in\{0,1\}$ (${\cal V}_{0}\cup{\cal V}_{1}={\cal V}$) 7:Extract watermark ${\mathbb{W}}=(w_{1},\ldots,w_{2nN})$ using ${\cal V}_{0}$ and ${\cal V}_{1}$ 8:for $i=1$ to $N$ do 9:     take $x_{i}$ 10:     while $x_{i}\neq(w_{j},\ldots,w_{j+n})$ do 11:         move $x_{i}$ along ${\mathbb{W}}$ bit by bit for $j=1,\ldots,n(2N-1)$ 12:     end while 13:     if $x_{i}=(w_{j},\ldots,w_{j+n})$ then 14:         ${\mathbb{P}}\leftarrow{\mathbb{P}}\cup(x_{i},(w_{j+n+1},\ldots,w_{j+2n}))$ 15:     end if 16:end for 17:return $\tilde{K}=(a_{0}\|a_{1})$ if the set ${\mathbb{P}}$ contains at least $\varepsilon\ell$ points on the straight line $f(x)=a_{0}+a_{1}x$
Algorithm 5 outlines the process of watermark extraction. The procedure starts by initializing the list $\mathbb{P}$ of points and computing the variables $x_{i}=x_{i}$ for $i=1,\dots,N$ using the LFSR. Two lists, ${\cal V}_{0}$ and ${\cal V}_{1}1$, are then created. Once these lists are prepared, the watermark ${\mathbb{W}}=(w_{1},\ldots,w_{2nN})$ is extracted. Each variable $x_{i}$ is processed sequentially, with the algorithm sliding $x_{i}$ bit by bit along ${\mathbb{W}}$ to search for a match. When a match is found, the next $n$ bits of ${\mathbb{W}}$ are assumed to be the suspected $f(x_{i})=(w_{j+n+1},\ldots,w_{j+2n})$, and the corresponding point is stored in the list $\mathbb{P}$. Processing each $x_{i}$ results in at most $2n(N-1)$ points, meaning that the set $\mathbb{P}$ has a cardinality of at most $2n(N-1)N$. If the text has been tampered with by an adversary, some points may no longer lie on the straight line $f(x)=a_{0}+a_{1}x$, and these outliers can be treated as random points. As long as enough points remain on the correct straight line, it can still be reliably determined, provided that adversarial distortions do not generate competing lines with a similar number of points. The primary challenge in this process is identifying the straight line containing the largest subset of points from the set, which is a well-known problem called the Maximum Collinear Points (MCP) problem.

### v2

@@ Abstract
The rapid advancement of LLMs (Large Language Models) has established them as a foundational technology for many AI- and ML-powered human–computer interactions. A critical challenge in this context is the attribution of LLM-generated text – either to the specific language model that produced it or to the individual user who embedded their identity via a so-called multi-bit watermark. This capability is essential for combating misinformation, fake news, misinterpretation, and plagiarism. One of the key techniques for addressing this challenge is digital watermarking.
This work presents a watermarking scheme for LLM-generated text based on Lagrange interpolation, enabling the recovery of a multi-bit author identity even when the text has been heavily redacted by an adversary. The core idea is to embed a continuous sequence of points $(x,f(x))$ that lie on a single straight line. The $x$-coordinates are computed pseudorandomly using a cryptographic hash function $H$ applied to the concatenation of the previous token’s identity and a secret key $s_{k}$. Crucially, the $x$-coordinates do not need to be embedded into the text – only the corresponding $f(x)$ values are embedded. During extraction, the algorithm recovers the original points along with many spurious ones, forming an instance of the Maximum Collinear Points (MCP) problem, which can be solved efficiently. Experimental results demonstrate that the proposed method is highly effective, allowing the recovery of the author’s identity even when as few as three genuine points remain after adversarial manipulation.
@@ 1 Introduction
The growing adoption of Large Language Models (LLMs) offers a wide array of practical benefits, including automated text correction, assistance with writing, and content refinement. However, alongside these advantages lies a more troubling aspect: the potential misuse of LLMs. One such concern arises when individuals present AI-generated content as their own, falsely claiming authorship. This challenge naturally raises an important question: how can we embed a watermark into LLM-generated text that allows a recipient to reliably trace its origin, whether to a specific user or to the LLM system itself? The objective of such watermarking is to make it possible to attribute the text to its true source, thereby establishing authorship beyond reasonable doubt. This capability is especially critical in combating misinformation, propaganda, automated plagiarism, and the spread of fake news.
@@ 2 Overview and Taxonomy of Watermarking in LLMs
– the output probability distribution (logits) is subtly perturbed before sampling tokens, embedding statistical patterns without altering the model weights. The work [12] splits the collection of tokens into green and red subsets. The green tokens encode watermarks while red ones mark fake ones (zero-bit watermarking). Wong et. al. [22] describes their logits-to-text watermarking, which jointly optimises encoder-decoder architecture for high-fidelity and high-detectability watermarks. The GumbelSoft watermarking described in [10] improves sampling diversity while maintaining robust detection by refining Gumbel-based logits control.
@@ 2.1 Contributions
    We design a watermarking scheme that embeds points from a straight line $f(x)$ into LLM-generated text. The $x$-coordinates are computed using a cryptographic hash function $H()$ applied to the identity of the preceding token and a secret key $s_{k}$. Notably, there is no need to embed these coordinates explicitly, as the verifier can reconstruct them directly from the text. Only the corresponding $f(x)$ values are embedded. To extract the watermark, the verifier must solve an instance of the Maximum Collinear Points (MCP) problem. We show that several algorithms exist for this task, with the most efficient ones achieving quadratic complexity, i.e., ${\cal O}(N^{2})$, where $N$ is the number of candidate points.
@@
    The proposed scheme can be naturally extended to support long multi-bit identities. We present two options: (1) using multiple lines, each encoding a distinct part of the identity, and (2) using higher-degree polynomials instead of lines. In the latter case, reconstructing the embedded polynomial corresponds to solving the Maximum Co-Polynomial Points problem. However, existing algorithms for this task have complexities that grow with the polynomial’s degree.
Table 2 compares some multi-bit watermarking schemes with the one presented in this work. The scheme by Qu et al. [18] proposes a robust multi-bit watermarking that encodes user identity by mapping segments of the message to selected token positions using pseudo-random assignment, reinforced with Reed–Solomon error correction. The paper [11] by Haoyu Jiang et al. proposes a multi-bit watermarking framework designed to credibly identify texts generated by LLMs. The core idea involves a trusted third party (TTP) coordinating with multiple LLM vendors to embed and later extract identity-carrying watermarks.
Table 1: Some LLM-generated text multi-bit watermarking schemes, where $b$ is the number of bits in user identity, $k$ is the number of segments, $T$ is the number of tokens in the text and $N$ is the number of points
Authors Tools Extraction Algorithm Complexity
Qu et al [18] RS Codes, Hash RS decoding ${\cal O}(2^{b})$ or ${\cal O}(k2^{\nicefrac{{b}}{{k}}})$
Jiang et al [11] TTP (digital signatures), Hash Deterministic mapping ${\cal O}(T+b)$
This work Lagrange Interpolation, Hash MCP ${\cal O}(N^{2})$
The rest of the paper is structured as follows. Section 3 focuses on the quality and security properties required from a watermarking scheme. Section 4 describes the tools and building blocks used in the design of our watermarking schemes. Section 5 introduces our basic watermarking scheme along with algorithms for recovering watermarking bits. Section 6 presents probabilistic analysis evaluating the likelihood that an adversary can implant a fake point that would be accepted as genuine. Section 7 details our experimental results. Section 8 demonstrates how our watermarking scheme can be extended to support long multi-bit identities. Finally, Section 9 concludes the paper.
@@ 3 Quality and Security Requirements for LLM-text Watermarking
In this work, we focus on a specific security goal: enabling the owner of an LLM-generated text to prove that the text contains a substantial portion of their original content, from which a multi-bit identity $K$ can be reliably recovered. This identity serves as a cryptographic witness of authorship or ownership. Our approach is grounded in the use of Lagrange interpolation over points in a two-dimensional plane. Each point encodes a fragment of the identity. To ensure unpredictability and prevent pattern leakage, we employ a cryptographic hash $H()$ to determine the $x$-coordinates of the interpolation points.
@@ 5 Watermark based on Lagrange Interpolation
In many real-world scenarios, the owner of an LLM-generated text may wish to assert authorship by demonstrating that the embedded watermark allows recovery of a multi-bit identity $K$. For this to be effective, the extraction algorithm must be able to retrieve the identity even if the text has been altered or partially redacted by an adversary. To support robust and verifiable authorship attribution, any sufficiently long fragment of the watermarked text should carry enough information to reconstruct $K$, serving as undeniable proof of origin.
To achieve this, we propose using Lagrange interpolation to recover a secret polynomial $f(x)$ whose coefficients encode the identity $K$. Each watermarking point is of the form $(x,f(x))$, where the $x$-coordinates are generated pseudorandomly using a cryptographic hashing. This design ensures that even if only a fraction of these points survive tampering, they suffice to recover $K$. The scheme is naturally resilient to partial modifications, making it suitable for adversarial environments where text may be edited with the intent to obscure provenance.
As previously mentioned, we focus on the simplest nontrivial case where $f(x)$ is a degree-1 polynomial, i.e., a straight line of the form $f(x)=a_{0}+a_{1}x\in GF(2^{n})[x]$. The secret identity is encoded as $K=(a_{0}\|a_{1})\in GF^{2}(2^{n})$. A straightforward strategy would embed $N$ watermark points directly onto this line, which requires a text of at least $2nN$ tokens. To improve efficiency, we propose a dense embedding scheme where each value $y_{i}=f(x_{i})$ is embedded using $n$ tokens, and the $x$-coordinate $x_{i}$ is derived from the identity of the preceding token. Formally, given a sequence of tokens $S_{i-1},S_{i},\ldots,S_{i+n}$, we define:
$x_{i}=H(ID_{S_{i-1}},s_{k}),$
and use the subsequent block $(S_{i},\ldots,S_{i+n})$ to encode $f(x_{i})$. This construction requires only $nN+1$ tokens to embed $N$ points, reducing overhead compared to the naive $2nN$-token requirement. The first token $S_{0}$ is not watermarked; it serves solely to generate the initial $x$-coordinate.
Algorithm 5 outlines the watermark embedding process based on Lagrange interpolation. Our approach diverges from common designs where the token vocabulary is repartitioned for each embedding step. While dynamically splitting ${\cal V}_{0}$ and ${\cal V}_{1}$ per token can enhance security, it introduces unnecessary computational overhead, particularly on the verifier’s side.
To embed a watermark bit $i\in\{0,1\}$, the LLM is queried to produce token logits. The probabilities of top-ranked tokens are then adjusted so that tokens from the appropriate subset ${\cal V}_{i}$ are favored subject to a parity constraint. This constraint can be defined either as the XOR of token ID bits or as the parity of the output of a cryptographic hash function $H(ID_{\text{token}},s_{k})$ (see Section 4.2). In this construction, the sets ${\cal V}_{0}$ and ${\cal V}_{1}$ are kept static across the embedding process, ensuring efficient and consistent watermark detection.
@@
Algorithm 1: Watermark Embedding Secret Identity $K$ 1:Input: Autoregressive language model LLM used to generate a text of length $T=nN+1$, prompt sequence $PR$, token set ${\cal V}$, $n$-bit secret key $s_{k}$, cryptographic hash $H()$ and straight line $f(x)=a_{0}+a_{1}x$ with the secret identity $K=(a_{0}\|a_{1})$, where computations are done in $GF(2^{n})$. 2:Output: Public LLM text ${\cal T}=S_{0,n},S_{1,1},\dots,S_{N,n}$.   3:Using secret key $s_{k}$, create two lists ${\cal V}_{0}$, ${\cal V}_{1}$, where ${\cal V}_{i}$ are tokens carrying watermark $i\in\{0,1\}$ (${\cal V}_{0}\cup{\cal V}_{1}={\cal V}$) 4:Initialize ${\cal T}=S_{0,n}\leftarrow LLM(PR)$ 5:for $i=1$ to $N$ do 6:     $x_{i}=(x_{i,1},x_{i,2},\ldots,x_{i,n})\leftarrow H(ID_{S_{i-1,n}},s_{k})$triangleright $x$- coordinate 7:     $y_{i}=f(x_{i})=(y_{i,1},y_{i,2},\ldots,y_{i,n})$triangleright value $y=f(x)$ 8:     for $j=1$ to $n$ do 9:         Compute logits: $v=\text{LLM}(PR,S_{0,n},S_{1,1},\dots,S_{i,j-1})$ 10:         Bias the logits: 11:         if $y_{i,j}=0$ then add bias to $\hat{v}[w]=v+\delta$ for all $w\in{\cal V}_{0}$ 12:         else add bias to $\hat{v}[w]=v+\delta$ for all $w\in{\cal V}_{1}$ 13:         end if 14:         Append $S_{i,j}$ to $\cal T$ by sampling from biased logits $\hat{v}$ 15:     end for 16:end for 17:return $\cal T$
@@
The algorithm processes text block by block. For each block (lines 6,7), two binary sequences are prepared. The first is a block generated by the hash, which consists of an $x$-coordinate. The second block is $y_{i}=f(x_{i})$. The loop from line 8 to 15 injects bits $y_{i}$ into the LLM text.
@@
Algorithm 2: Watermark Extraction 1:Input: LLM text $\widetilde{\cal T}=(\tilde{S}_{0},\tilde{S}_{1},\dots,\tilde{S}_{nN})$, token set ${\cal V}$, $n$-bit secret key $s_{k}$, $n$-bit cryptographic hash $H()$, which generates $x$-coordinates of unknown $f(x)$, where points $(x_{i},f(x_{i}))$ are embedded into the text according to Algorithm 5; $i=1,\ldots,N$, computations are done in $GF(2^{n})$. 2:Output: The secret user identity $K=(a_{0}\|a_{1})$   3:Create an empty list of points ${\mathbb{P}}=\varnothing$ of $(x,f(x))$ 4:Using secret key $s_{k}$, create two lists ${\cal V}_{0}$, ${\cal V}_{1}$, where ${\cal V}_{i}$ are tokens carrying watermark $i\in\{0,1\}$ (${\cal V}_{0}\cup{\cal V}_{1}={\cal V}$) 5:Extract watermark ${\mathbb{W}}=(w_{1},\ldots,w_{nN})$ using ${\cal V}_{0}$ and ${\cal V}_{1}$ 6:for $i=0$ to $n(N-1)$ do 7:     $x_{i}=(x_{i,1},x_{i,2},\ldots,x_{i,n})\leftarrow H(ID_{\tilde{S}_{i}},s_{k})$ 8:     $y_{i}=(w_{i+1},\ldots,w_{i+1+n})$ 9:     ${\mathbb{P}}\leftarrow{\mathbb{P}}\cup{(x_{i},y_{i})}$ 10:end for 11:Solve the maximum collinear points instance for ${\mathbb{P}}$ and find the line $f(x)=a_{0}+a_{1}x$, which contains them 12:return $\tilde{K}=(a_{0}\|a_{1})$
Algorithm 5 outlines the process of watermark extraction. The procedure begins by initializing the list $\mathbb{P}$ of candidate points. Two token subsets, ${\cal V}_{0}$ and ${\cal V}_{1}$, are then constructed. Once these are prepared, the watermark sequence ${\mathbb{W}}=(w_{1},\ldots,w_{nN})$ is extracted from the modified text. The watermark bits and their corresponding tokens are processed sequentially. For each bit, the algorithm computes the coordinate $x_{i}=H(ID_{\tilde{S}_{i-1}},s_{k})$ using the identity of the preceding token. It then assigns the value $y_{i}$ as the sequence $(w_{i+1},\ldots,w_{i+1+n})$, and the resulting point $(x_{i},y_{i})$ is appended to the list $\mathbb{P}$.
As a result, the extraction algorithm builds a list of $N$ genuine points that lie on the original line $f(x)$, along with $n(N-1)+1$ additional points that are scattered across the 2D plane. An adversary may both insert fake points and remove genuine ones, attempting to obscure the original structure. Nonetheless, as long as a sufficient number of authentic points remain on the correct straight line, the identity $K$ can still be reliably recovered – assuming the adversary does not manage to create a competing line with a comparable number of points. The key computational challenge becomes identifying the straight line that contains the largest subset of points from the entire set – a classic geometric problem known as the Maximum Collinear Points (MCP) problem.